OpenStudy (faman39):

Rick tipped the pizza delivery driver $3 for pizzas that cost $18. Approximately what percent did he tip the driver? Round your answer to the nearest percent.

OpenStudy (wasiqss):

(3/18)*100= 17% approximate

OpenStudy (callisto):

amount he tipped / the cost *100%

Directrix (directrix):

$24 pizza $3 tip so 3/24 times !00 = percentage tip --> 12.5 %

OpenStudy (wasiqss):

directrix you love posting wrong answers ?:D :D

OpenStudy (anonymous):

(3/18)*100 =300/18 =16.7% =17%